OpenStudy (anonymous):

please teach me how to solve a-b=2 a+b=2

OpenStudy (anonymous):

you could set \[a=b+2\] then substitute it for a in the other equation.

OpenStudy (anonymous):

Or you can just add the two equations.

OpenStudy (anonymous):

a-b = 2 + a+b = 2 ======= 2a + 0 = 4

OpenStudy (anonymous):

Then you can find a, and plug into one of the equations to solve for b.